How to find the equation of trajectory using conservation of energy theorem


1
Expert's answer
2021-01-04T14:40:07-0500

"E_p+Ek = const"

"E_p,E_k-\text{potential and kinetic energy of the body}"

"E_p = mgh"

"Ек = \frac{mV^2}{2}"

"\text{let }E_k =0;E_p=mgh_{max};const= mgh_{max}"

"mgh +\frac{mV^2}{2}=mgh_{max}"

"h= h_{max}-\frac{V^2}{2g}"

"h(t)= h_{max}-\frac{V(t)^2}{2g};V(t) -\text{change in speed over time}"

Answer:"h(t)=h_{max}-\frac{V^2(t)}{2g};V(t)-\text{changein speed over time}"